### v1

# Towards Unified Latent Space for 3D Molecular Latent Diffusion Modeling

Abstract
3D molecule generation is crucial for drug discovery and material science, requiring models to process complex multi-modalities, including atom types, chemical bonds, and 3D coordinates. A key challenge is integrating these modalities of different shapes while maintaining SE(3) equivariance for 3D coordinates. To achieve this, existing approaches typically maintain separate latent spaces for invariant and equivariant modalities, reducing efficiency in both training and sampling. In this work, we propose Unified Variational Auto-Encoder for 3D Molecular Latent Diffusion Modeling (UAE-3D), a multi-modal VAE that compresses 3D molecules into latent sequences from a unified latent space, while maintaining near-zero reconstruction error. This unified latent space eliminates the complexities of handling multi-modality and equivariance when performing latent diffusion modeling. We demonstrate this by employing the Diffusion Transformer–a general-purpose diffusion model without any molecular inductive bias–for latent generation. Extensive experiments on GEOM-Drugs and QM9 datasets demonstrate that our method significantly establishes new benchmarks in both de novo and conditional 3D molecule generation, achieving leading efficiency and quality.
Towards Unified Latent Space for 3D Molecular Latent Diffusion Modeling
Yanchen Luo1, Zhiyuan Liu2, Yi Zhao1, Sihang Li1, Kenji Kawaguchi2, Tat-Seng Chua2, Xiang Wang1 1University of Science and Technology of China 2National University of Singapore luoyanchen@mail.ustc.edu.cn
1 Introduction
The discovery of novel molecules is fundamental to advancements across various scientific fields, particularly in drug and material development. Given significant progress has been made in designing 2D molecular graphs Jin et al. (2018); Irwin et al. (2022); Polykovskiy et al. (2020); Shi et al. (2020), recent research has increasingly focused on the generation of 3D molecules Hoogeboom et al. (2022); Huang et al. (2023b). Unlike 2D molecular generation, which focuses on forming valid molecular structures based on chemical bonds, 3D generation must also predict 3D atomic coordinates that align with the 2D structure. Accurate 3D molecule generation is essential to power many important applications, such as structure-based drug design Zhang et al. (2023) and inverse molecular design targeting quantum chemical properties Bao et al. (2023).
Figure 1: (a) A 3D molecule is characterized by features of distinct modalities. (b) Conventional approaches use separate latent spaces for equivariant (3D coordinates) and invariant (2D features) components, leading to fragmented molecular representations. (c) Our UAE-3D establishes a unified SE(3)-equivariant latent space that jointly encodes both modalities
Table 1: Comparing the reconstruction error between UAE-3D and GeoLDM’s VAE. Atom/Bond accuracy measures correct atom/bond type predictions. Coordinate RMSD quantifies geometric deviation from the ground truth 3D coordinates.
Metric GeoLDM Xu et al. (2023) UAE-3D (Ours)
Atom Accuracy (%) 98.6 100.0
Bond Accuracy (%) 96.2 100.0
Coordinate RMSD (Å) 0.1830 0.0002
3D molecule generation is challenging due to its multi-modal nature. As shown in Figure 1(a), a 3D molecule consists of features of three distinct modalities: atom types, atomic coordinates, and chemical bonds. This requires the generation model to handle both discrete (e.g., atom types) and continuous features (e.g., coordinates), while also addressing differences in feature shapes (atom-wise v.s. edge-wise features). Worse still, the modality of 3D coordinate requires special care to respect rotational and translational equivariance, whereas other modalities do not, further complicating the task. Mitigating this multi-modal challenge, prior works mostly process each modality in separate latent spaces: some maintain separate diffusion processes for each modality Huang et al. (2023b); Vignac et al. (2023); Le et al. (2024); Hua et al. (2023), while others interleave the prediction of different modalities across autoregressive generation steps Gebauer et al. (2019, 2021). However, we argue that handling each modality separately complicates the model design, reducing both the training and sampling efficiency. Moreover, processing each modality separately risks compromising the consistency between them. These issues raise a critical question: can we design a unified generative model that seamlessly integrates all three modalities of 3D molecule generation?
To answer this research question, we propose to build a multi-modal latent diffusion model (LDM) Rombach et al. (2022) for the unified generative modeling of 3D molecules. LDM extends the diffusion paradigm Ho et al. (2020) by operating in a compressed latent space learned through variational autoencoders (VAEs) Kingma and Welling (2014), offering improved computational efficiency and generation quality. For 3D molecules, we can build a multi-modal VAE that compresses all three modalities in a single unified latent space. Scrutinizing the previous works, we identify the following challenges to achieve this purpose:
  • Building a Single Unified Latent Space for 3D Molecules. While previous studies have explored LDMs for 3D molecule and protein generation Xu et al. (2023); Kong et al. (2024), they fail to build a unified latent space that integrates all modalities. Their difficulty arises from the reliance on neural networks with baked-in 3D equivariance Satorras et al. (2021); Huang et al. (2023b), which mostly maintain separate latent spaces for 3D equivariant and invariant modalities (cf. Figure 1(b)).
  • Ensuring Near-Lossless Compression. Maintaining a low reconstruction error for the multi-modal VAE is critical, as even small errors can result in invalid or unstable 3D structures. Furthermore, imprecise latents that cannot reproduce the original molecule can propagate their errors to the LDM, easily disrupting the generative modeling. Despite its critical importance, reconstruction error has been largely overlooked in previous works Xu et al. (2023); Kong et al. (2024) (cf. Table 1).
In this work, we introduce Unified Variational Auto-Encoder for 3D Molecular Latent Diffusion Modeling (UAE-3D), a VAE that can compress the multi-modal features of 3D molecules into a unified latent space while maintaining near zero (100% atom/bond accuracy and 2E-4 coordinate RMSD) reconstruction error. To obtain a latent space integrating both 3D equivariant and invariant modalities, we draw inspiration from the “bitter lesson” Wang et al. (2024) of 3D molecules: rather than using models with intricate, baked-in 3D equivariance, UAE-3D trains a neural network to “learn” 3D equivariance through our tailor-made SE(3) augmentations, encouraging the transformation on the input coordinates are reflected equivariantly on the output coordinates. Moreover, UAE-3D employs the Relational Transformer Diao and Loynd (2023) as its encoder, leveraging its scalability and flexibility to incorporate both atom-wise and edge-wise features. A Transformer-based Vaswani et al. (2017) decoder is jointly trained with the encoder to reconstruct both 3D invariant and equivariant molecular features.
Despite its simplicity, UAE-3D can compress 3D molecules as token sequences in a unified latent space, eliminating the complexities of handling multi-modalities and 3D equivariance in latent diffusion modeling. As shown in Table 1, UAE-3D reaches lossless reconstruction of atom and bond types, with a near-zero coordinate RMSD. To further demonstrate its effectiveness, we employ the Diffusion Transformer (DiT) Peebles and Xie (2023), a general-purpose diffusion backbone without any molecular inductive bias, to model UAE-3D’s latents. We show that DiT can successfully generate stable and valid 3D molecules, with significantly improved training (by 2.7 times) and sampling efficiency (by 7.3 times). We refer to this LDM pipeline as UDM-3D: Unified Latent Diffusion Modeling for 3D Molecule Generation.
Extensive experiments demonstrate that UDM-3D, powered by UAE-3D, achieves state-of-the-art results in both de novo and conditional 3D molecule generation on the QM9 Ramakrishnan et al. (2014) and GEOM-Drugs Axelrod and Gómez-Bombarelli (2020) datasets. Ablation studies show the effectiveness of our key components.
Figure 2: Overview of the UDM-3D and UAE-3D models. The UAE-3D encodes 3D molecules from molecular space into a unified latent space, integrating multi-modal features such as atom types, chemical bonds, and 3D coordinates. This unified latent space is then utilized by the UDM-3D, which employs a DiT to perform generative modeling. Then the denoised latents are decoded back into 3D molecules.
2 Background: 3D Molecular Latent Diffusion Models
In this work, we explore latent diffusion models (LDMs) for 3D molecule generation. A 3D molecular LDM involves a 3D molecular VAE to compress 3D molecules into the latent space, where a diffusion model performs generative modeling. Below, we briefly introduce both components.
2.1 Molecular Variational Auto-Encoding
Notations. A 3D molecule is represented by $\mathbf{G}=\langle\mathbf{F},\mathbf{E},\mathbf{X}\rangle$, where $\mathbf{F}\in\mathbb{R}^{|\mathcal{V}|\times d_{1}}$ is the atom feature matrix (e.g., atomic numbers), $\mathbf{E}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|\times d_{2}}$ is the bond feature matrix (e.g., bond connectivity and type), and $\mathbf{X}\in\mathbb{R}^{|\mathcal{V}|\times 3}$ is the 3D atom coordinate matrix. $d_{1}$ and $d_{2}$ are dimensions of atom and bond features. $\mathcal{V}$ is $\mathbf{G}^{\prime}s$ set of atoms and $|\mathcal{V}|$ denotes the number of atoms.
A molecular VAE consists of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. Given a 3D molecule $\mathbf{G}$, the encoder $\mathcal{E}$ maps it into a sequence of latent tokens $\mathcal{E}(\mathbf{G})=\mathbf{Z}=\{\boldsymbol{z}_{i}\in\mathbb{R}^{d}|i\in\mathcal{V}\}$. Each latent $\boldsymbol{z}_{i}$ is sampled from a Gaussian distribution $\mathcal{N}(\boldsymbol{z}_{i};\boldsymbol{\mu}_{i},\boldsymbol{\sigma}_{i})$ using the reparameterization trick Kingma and Welling (2014), where $\boldsymbol{\mu}_{i}$ and $\boldsymbol{\sigma}_{i}$ are learned parameters for atom $i$. The decoder $\mathcal{D}$ then reconstructs $\mathbf{G}$ from these latents: $\hat{\mathbf{G}}=\mathcal{D}(\mathbf{Z})$. The VAE is trained to minimize a reconstruction loss and a regularization term to ensure the latent space follows a standard Gaussian prior:
$\mathcal{L}_{\operatorname{VAE}}=$ (1)
$\mathbb{E}_{\mathbf{G}\sim p_{\operatorname{data}}}\left[\|\mathbf{G}-\hat{\mathbf{G}}\|+D_{\operatorname{KL}}(q(\mathbf{Z}|\mathbf{G})\|p(\boldsymbol{z}))\right],$
where $D_{\operatorname{KL}}$ denotes the Kullback-Leibler divergence; $q(\mathbf{Z}|\mathbf{G})=\prod_{i\in\mathcal{V}}\mathcal{N}(\mathbf{z}_{i};\boldsymbol{\mu}_{i},\boldsymbol{\sigma}_{i})$ is the approximated posterior distribution; and $\|\mathbf{G}-\hat{\mathbf{G}}\|$ measures the difference between the original and the predicted graph, whose definition varies across different works. In Xu et al. (2023), $\|\mathbf{G}-\hat{\mathbf{G}}\|$ combines an MSE loss for 3D coordinates and a cross-entropy loss for atom types.
2.2 Diffusion Model
Building on the VAE’s latent $\mathbf{Z}$ for 3D molecules, a molecular LDM performs generative modeling with a diffusion model Rombach et al. (2022); Kingma et al. (2021). In the forward diffusion process, we gradually add noise to the original latent $\mathbf{Z}^{(0)}=\mathbf{Z}$ following:
$\mathbf{Z}^{(t)}\sim\mathcal{N}(\mathbf{Z}^{(t)};\sqrt{\bar{\alpha}^{(t)}}\mathbf{Z}^{(0)},(1-\bar{\alpha}^{(t)})\mathbf{I}),$ (2)
where $t\in(0,1]$ is the diffusion timestep, and $\bar{\alpha}^{(t)}$ is a hyperparameter controlling the signal and noise ratio. At the final step ($t=1$), $\mathbf{Z}^{(1)}$ approaches the prior distribution, an isotropic Gaussian distribution: $\mathbf{Z}^{(1)}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I})$. In practice, $\mathbf{Z}^{(t)}$ is sampled as $\mathbf{Z}^{(t)}=\sqrt{\bar{\alpha}^{(t)}}\mathbf{Z}^{(0)}+\sqrt{1-\bar{\alpha}^{(t)}}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I})$.
Given the noised latents $\mathbf{Z}^{(t)}$, a diffusion model $\boldsymbol{\epsilon}_{\theta}(\mathbf{Z}^{(t)},t)$ trained to predict the added noise $\boldsymbol{\epsilon}$ by minimizing the MSE loss $\|\epsilon_{\theta}(\mathbf{Z}^{(t)},t)-\boldsymbol{\epsilon}\|^{2}$. Once trained, new latents $\hat{\mathbf{Z}}^{(0)}$ are sampled using the diffusion model $\boldsymbol{\epsilon}_{\theta}$ via the iterative denoising process Ho et al. (2020), and the decoder $\mathcal{D}$ reconstructs the corresponding 3D molecule via $\hat{\mathbf{G}}=\mathcal{D}(\hat{\mathbf{Z}}^{(0)})$.
[…]
3 Methodology
In this section, we propose the UAE-3D, a multi-modal variational auto-encoder designed to effectively compress the diverse modalities of 3D molecules into a unified latent space. Based on UAE-3D, we introduce UDM-3D, an LDM for 3D molecule generation.
3.1 UAE-3D: Unified Variational Auto-Encoder for 3D Molecular Latent Diffusion Modeling
UAE-3D is designed to address the complexities of multi-modal and equivariance of molecular data by compressing the atomic features, bond features, and atomic coordinates into a unified latent space. This is achieved through three key components: (1) a Relatinoal-Transformer Diao and Loynd (2023) that effectively integrates the multi-modal features into token sequences, (2) a Transformer-based Vaswani et al. (2017) decoder and the reconstruction loss for 3D molecule reconstruction; and (3) our tailor-made data augmentations to train the model to learn 3D equivariance.
Compressing 3D Molecules with Relational Transformer (R-Trans).
Given a 3D molecular $\mathbf{G}=\langle\mathbf{F},\mathbf{E},\mathbf{X}\rangle$, we compute its initial atom-wise and edge-wise embeddings as:
$\mathbf{H}^{n}$ $=\operatorname{MLP}([\mathbf{X};\mathbf{F}])$ $\in\mathbb{R}^{|\mathcal{V}|\times d},$ (3)
$\mathbf{H}^{e}$ $=\operatorname{MLP}([\mathbf{E};\mathbf{D}])$ $\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|\times d},$ (4)
$\mathbf{D}_{ij}$ $=\operatorname{GBF}(\|\mathbf{X}_{i}-\mathbf{X}_{j}\|^{2})$ $\in\mathbb{R}^{d},$ (5)
where $\operatorname{GBF}(\cdot)$ implements the Gaussian basis functions to expand inter-atomic distances as feature vectors Thölke and Fabritiis (2022); $d$ is the embedding dimension. The initial node embeddings $\mathbf{H}^{n}$ combines both atomic features and its 3D coordinates, while the edge embeddings $\mathbf{H}^{e}$ incorporate bond features and inter-atomic distances. $\mathbf{H}^{n}$ and $\mathbf{H}^{e}$ together effectively represent all 3D molecular information for the subsequent encoding. After obtaining the initial embeddings, we process them through $L$ layers of R-Trans blocks Diao and Loynd (2023):
$\tilde{\mathbf{H}}^{n}=\operatorname{R-Trans}(\mathbf{H}^{n},\mathbf{H}^{e})\in\mathbb{R}^{|\mathcal{V}|\times d},$ (6)
where $\tilde{\mathbf{H}}^{n}$ denotes the updated atom embedding for the next layer, computed as follows:
$\mathbf{Q}_{ij}=[\mathbf{H}^{n}_{i};\mathbf{H}^{e}_{ij}]\mathbf{W}^{q},$ $\mathbf{Q}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|\times d},$ (7)
$[\mathbf{K}_{ij};\mathbf{V}_{ij}]=[\mathbf{H}^{n}_{j};\mathbf{H}^{e}_{ij}]\mathbf{W}^{kv},$ $\mathbf{K},\mathbf{V}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|\times d},$ (8)
$\boldsymbol{\alpha}_{ij}=\text{softmax}_{j}\left(\frac{\mathbf{Q}_{ij}\mathbf{K}_{ij}^{\top}}{\sqrt{d}}\right),$ $\boldsymbol{\alpha}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|},$ (9)
$\hat{\mathbf{H}}_{i}^{n}=\sum_{j}\boldsymbol{\alpha}_{ij}\mathbf{V}_{ij},$ $\hat{\mathbf{H}}^{n}\in\mathbb{R}^{|\mathcal{V}|\times d},$ (10)
$\tilde{\mathbf{H}}_{i}^{n}=\operatorname{MLP}(\hat{\mathbf{H}}_{i}^{n}),$ $\tilde{\mathbf{H}}^{n}\in\mathbb{R}^{|\mathcal{V}|\times d},$ (11)
where $\mathbf{W}^{q}$ and $\mathbf{W}^{kv}$ are linear projectors. Unlike the original R-Trans, our implementation does not update the edge embedding $\mathbf{H}^{e}$ across layers. We find that this method is more efficient without compromising performance.
Compared to Transformer Vaswani et al. (2017), R-Trans can effectively integrate edge embeddings $\mathbf{H}^{e}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|\times d}$ and atom embeddings $\mathbf{H}^{n}\in\mathbb{R}^{|\mathcal{V}|\times d}$, despite their different shapes, making it well-suited for 3D molecule encoding. This integration occurs during the computation of queries, keys and values (Equation (7) and (8)), ensuring the output atom embeddings fully incorporate edge information.
Decoder. We employ a standard Transformer Vaswani et al. (2017) as decoder, with three additional predictors to reconstruct complete 3D molecules: atom types, bond types, and atom coordinates. Unlike the encoder, our decoder includes no molecule-specialized design, because it processes sequences of latents. Given the encoder’s latents $\mathbf{Z}$, our Transformer decoder provides the representation:
$\mathbf{P}=\operatorname{Transformer}(\mathbf{Z})\in\mathbb{R}^{|\mathcal{V}|\times d}.$ (12)
We then obtain predictions for three molecular modalities:
$\hat{\mathbf{X}}_{i}$ $=\operatorname{MLP}_{1}(\mathbf{P}_{i})$ $\in\mathbb{R}^{3},$ (13)
$\hat{\mathbf{F}}_{i}$ $=\operatorname{MLP}_{2}(\mathbf{P}_{i})$ $\in\mathbb{R}^{N_{a}},$ (14)
$\hat{\mathbf{E}}_{ij}$ $=\operatorname{MLP}_{3}(\mathbf{P}_{i}+\mathbf{P}_{j})$ $\in\mathbb{R}^{N_{b}},$ (15)
where $N_{a}$ and $N_{b}$ are the number of atom and bond types.
Reconstruction Loss. We define the distance between our reconstructed graph and the original graph $\|\hat{\mathbf{G}}-\mathbf{G}\|$, which includes multiple components to ensure the fidelity of the generated 3D molecules:
$\mathcal{L}_{\text{atom}}$ $=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\operatorname{CrossEntropy}(\hat{\mathbf{F}}_{i},\mathbf{F}_{i}),$ (16)
$\mathcal{L}_{\text{coordinate}}$ $=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\operatorname{MSE}(\hat{\mathbf{X}}_{i},\mathbf{X}_{i}),$ (17)
$\mathcal{L}_{\text{bond}}$ $=\frac{1}{|\mathcal{V}|^{2}}\sum_{i,j\in\mathcal{V}}\operatorname{CrossEntropy}(\hat{\mathbf{E}}_{ij},\mathbf{E}_{ij}),$ (18)
$\mathcal{L}_{\text{distance}}$ $=\frac{1}{|\mathcal{V}|^{2}}\sum_{i,j\in\mathcal{V}}\operatorname{MSE}(\|\hat{\mathbf{X}}_{i}-\hat{\mathbf{X}}_{j}\|,\mathbf{D}_{ij}),$ (19)
where $\mathcal{L}_{\text{atom}}$ and $\mathcal{L}_{\text{bond}}$ are cross-entropy losses for atom and bond types; $\mathcal{L}_{\text{coordinate}}$ measures the MSE between the predicted and the ground truth coordinates; $\mathcal{L}_{\text{distance}}$ is an additionally constraint on the predicted coordinates, to ensure correct inter-atomic distances. The overall reconstruction loss is a weighted sum of these components:
$\mathcal{L}_{\text{recon}}=\boldsymbol{\gamma}\cdot[\mathcal{L}_{\text{atom}},\mathcal{L}_{\text{bond}},\mathcal{L}_{\text{coordinate}},\mathcal{L}_{\text{distance}}]^{\top},$ (20)
where $\boldsymbol{\gamma}\in\mathbb{R}^{4}$ is a hyperparameter vector balancing the reconstruction terms. The final loss for UAE-3D combines $\mathcal{L}_{\text{recon}}$ with a KL regularization term:
$\mathcal{L}_{\text{UAE-3D}}=\mathcal{L}_{\text{recon}}+\beta\cdot\mathbb{E}_{\mathbf{G}\sim p_{\text{data}}}\left[D_{\operatorname{KL}}(q(\mathbf{Z}|\mathbf{G})\|p(\mathbf{Z}))\right],$ (21)
where the hyperparameter $\beta\in\mathbb{R}$ controls the KL regularization strength.
SE(3)-Equivariant Augmentations. To ensure the VAE follows SE(3)-equivariance, we apply $\text{SE}(3)$ transformations $\mathbf{R}\in\text{SE}(3)$ as data augmentation on the molecular 3D coordinates $\mathbf{X}$ Abramson et al. (2024). The VAE $\mathcal{D}(\mathcal{E}(\cdot))$ is trained to be equivariant to these transformations:
$\mathcal{L}_{\text{recon}}=\|\mathcal{D}(\mathcal{E}(\mathbf{R}\circ\mathbf{G}))-\mathbf{R}\circ\mathbf{G}\|,$ (22)
where $\mathbf{R}\circ\mathbf{G}=\langle\mathbf{F},\mathbf{E},\mathbf{R}(\mathbf{X})\rangle$ denotes applying the transformation $\mathbf{R}$ on the 3D coordinates of $\mathbf{G}$. In practice, $\mathbf{R}$ combines a random rotation from $\text{SO}(3)$ and a random translation from $\mathcal{N}(\mathbf{0},0.01\mathbf{I}_{3})$, where $0.01$ controls the range of the translation augmentation. When these SE(3) augmentations are applied, Equation (22) becomes equivalent to Equation (20), training the VAE to follow SE(3)-equivariance.
3.2 UDM-3D: Unified Latent Diffusion Modeling for 3D Molecule Generation
Diffusion Transformer (DiT). Given the unified latent space provided by our UAE-3D model, we now apply the DiT Peebles and Xie (2023) as our diffusion model $\boldsymbol{\epsilon}_{\theta}$ for 3D molecular latent generation. DiT’s promising performance in image generation underscores its superior capability in modeling latent sequences. Unlike a standard Transformer Vaswani et al. (2017), DiT replaces the layernorm Ba et al. (2016) by adaptive layernorm (adaLN) Perez et al. (2018), where the shift $\mathbf{y}$ and scale $\mathbf{b}$ parameters are dynamically generated based on the diffusion timestep $t$: $[\mathbf{y};\mathbf{b}]=\text{MLP}(\text{Embed}_{t}(t))$, where $\text{Embed}(\cdot)$ embeds timestep $t$ into a vector. These shift $\mathbf{y}$ and scale $\mathbf{b}$ parameters act as “gates”, selectively activating different dimension given different timesteps, enabling effective dynamic-scale denoising in diffusion.
Conditional Generation with Classifier-Free Guidance. For conditional generation, we directly combine the condition embedding $\text{Embed}_{c}(\mathbf{c})$ with the time embedding $\text{Embed}_{t}(\mathbf{t})$ when generating parameters for the adaLN:
$[\mathbf{y};\mathbf{b}]=\text{MLP}(\text{Embed}_{t}(t)+\text{Embed}_{c}(c)).$ (23)
In this way, we effectively inform the diffusion model $\boldsymbol{\epsilon}_{\theta}(\mathbf{Z}^{(t)},\mathbf{t},\mathbf{c})$ with the condition information.
To further enforce conditioning during inference, we employ the classifier-free guidance Ho and Salimans (2022) (CFG) to find $\mathbf{Z}$ that has a high $\log p(\mathbf{c}|\mathbf{Z})$. By Bayes’ rule, the gradient of this objective is $\nabla_{\mathbf{Z}}\log p(\mathbf{c}|\mathbf{Z})\propto\nabla_{\mathbf{Z}}\log p(\mathbf{Z}|\mathbf{c})-\nabla_{\mathbf{Z}}\log p(\mathbf{Z})$. Following Song et al. (2021), we can interprete the diffusion model $\boldsymbol{\epsilon}_{\theta}$’s output as score functions, and therefore maximize $p(\mathbf{c}|\mathbf{Z})$ using the modified denoising function:
$\tilde{\boldsymbol{\epsilon}}_{\theta}(\mathbf{Z}^{(t)},\mathbf{t},\mathbf{c})=(1+w)\boldsymbol{\epsilon}_{\theta}(\mathbf{Z}^{(t)},\mathbf{t},\mathbf{c})-w\boldsymbol{\epsilon}_{\theta}(\mathbf{Z}^{(t)},\mathbf{t}),$ (24)
where $w\in[0,+\infty)$ is a hyperparameter controlling the guidance strength; $\boldsymbol{\epsilon}_{\theta}(\mathbf{Z}^{(t)},\mathbf{t},\mathbf{c})$ is the conditional variant of $\boldsymbol{\epsilon}_{\theta}$, incorporating the property $\mathbf{c}$; $\boldsymbol{\epsilon}_{\theta}(\mathbf{Z}^{(t)},\mathbf{t})$ is the unconditioned variant that ignores $\mathbf{c}$. To train $\boldsymbol{\epsilon}_{\theta}$, we randomly drop condition $\mathbf{c}$ with a certain probability, allowing $\boldsymbol{\epsilon}_{\theta}$ to learn both conditional and unconditional distributions.
Table 2: Performance of de novo 3D molecule generation on GEOM-Drugs. * indicates results reproduced using official source codes, while other baseline results are taken from Huang et al. (2023b).
[…]
Train 1.000 1.000 1.000 0.000 0.585 0.999 0.584 00.251
CDGS 0.991 0.285 0.285 0.285 0.262 0.789 0.022 22.051
JODO 1.000 0.874 0.905 0.902 0.417 0.993 0.483 02.523
MiDi* 0.968 0.633 0.654 0.652 0.392 0.951 0.196 07.054
UDM-3D, ours 1.000 0.879 0.913 0.958 0.525 0.989 0.540 00.692
${}_{\text{3D}}$$\downarrow$ Bond $\downarrow$ Bond $\downarrow$ Dihedral $\downarrow$
Train 0.861 13.73 1.56E-04 1.81E-04 1.56E-04
EDM 0.831 31.29 4.29E-01 4.96E-01 1.46E-02
JODO 0.845 19.99 8.49E-02 1.15E-02 6.68E-04
MiDi* 0.750 23.14 1.17E-01 9.57E-02 4.46E-03
GeoLDM 0.843 30.68 3.91E-01 4.22E-01 1.69E-02
UDM-3D, ours 0.852 17.36 9.89E-03 5.11E-03 1.74E-04
Table 3: Performance of de novo 3D molecule generation on QM9. * indicates results reproduced using official source codes, while other baseline results are taken from Huang et al. (2023b).
[…]
Train 0.999 0.989 0.989 0.000 0.490 0.992 0.946 0.063
CDGS 0.997 0.951 0.936 0.860* 0.493 0.973 0.784 0.798
JODO 0.999 0.990 0.960 0.780* 0.522 0.986 0.934 0.138
MiDi* 0.998 0.980 0.954 0.769 0.501 0.979 0.882 0.187
UDM-3D, ours 0.999 0.983 0.972 0.948 0.508 0.987 0.897 0.161
${}_{\text{3D}}$$\downarrow$ Bond $\downarrow$ Bond $\downarrow$ Dihedral $\downarrow$
Train 0.994 0.877 5.44E-04 4.65E-04 1.78E-04
EDM 0.986 1.285 1.30E-01 1.82E-02 6.64E-04
MDM 0.992 4.861 2.74E-01 6.60E-02 2.39E-02
JODO 0.992 0.885 1.48E-01 1.21E-02 6.29E-04
MiDi* 0.983 1.100 8.96E-01 2.08E-02 8.14E-04
GeoLDM 0.989 1.030 2.40E-01 1.00E-02 6.59E-04
UDM-3D, ours 0.993 0.881 7.04E-02 9.84E-03 3.47E-04
Table 4: Mean Absolute Error (MAE) for conditional 3D molecule generation on QM9. Lower values represent better performance.
Method $\mu\ (\textnormal{D})$ $\alpha\ (\textnormal{Bohr}^{3})$ $C_{v}\ \left(\frac{\textnormal{cal}}{\textnormal{mol}}\textnormal{K}\right)$ $\varepsilon_{\textnormal{HOMO}}\ (\textnormal{meV})$ $\varepsilon_{\textnormal{LUMO}}\ (\textnormal{meV})$ $\Delta\varepsilon\ (\textnormal{meV})$
U-Bound 1.613 8.98 6.879 645 1457 1464
EDM 1.123 2.78 1.065 371 0601 0671
EEGSDE 0.777 2.50 0.941 302 0447 0487
GeoLDM 1.108 2.37 1.025 340 0522 0587
JODO 0.628 1.42 0.581 226 0256 0335
UDM-3D, ours 0.603 1.76 0.553 216 0247 0313
L-Bound 0.043 0.09 0.040 039 0036 0065
4 Experiments
In this section, we present the experimental results of our proposed unified molecular latent space approach (UAE-3D & UDM-3D) on 3D molecule generation tasks. We comprehensively evaluate UDM-3D’s performance on de novo 3D molecule generation and conditional 3D molecule generation with targeted quantum properties. Our experiments demonstrate significant improvements over state-of-the-art methods while maintaining computational efficiency.
4.1 Experimental Setup
Datasets. Following previous studies Hoogeboom et al. (2022); Huang et al. (2023b); Vignac et al. (2023), we conduct experiments on the popular QM9 Ramakrishnan et al. (2014) and the GEOM-Drugs Axelrod and Gómez-Bombarelli (2020) datasets.
  • QM9 Ramakrishnan et al. (2014): Contains 130k small organic molecules (up to 9 heavy atoms, 29 total atoms) with quantum chemical properties. Following standard practice Hoogeboom et al. (2022); Huang et al. (2023b), we use 100K/18K/13K train/val/test splits.
Baselines. For de novo 3D molecule generation, we compare UDM-3D with CDGS Huang et al. (2023a), JODO Huang et al. (2023b), MiDi Vignac et al. (2023), G-SchNet Gebauer et al. (2019), G-SphereNet Luo and Ji (2022), EDM Hoogeboom et al. (2022), MDM Huang et al. (2023c), GeoLDM Xu et al. (2023). For conditional 3D molecule generation, we compare UDM-3D with EDM Hoogeboom et al. (2022), EEGSDE Bao et al. (2023), GeoLDM Xu et al. (2023), and JODO Huang et al. (2023b).
4.2 De Novo 3D Molecule Generation
Experimental Setting. Generating a 3D molecule involves generating the structural validity (atom/bond features) and accurate spatial arrangements (3D coordinates). Therefore we adapt the comprehensive metrics from Huang et al. (2023b); Hoogeboom et al. (2022):
  • 2D Metrics: Atom stability, validity & completeness (V&C), validity & uniqueness (V&U), validity & uniqueness & novelty (V&U&N), similarity to nearest neighbor (SNN), fragment similarity (Frag), scaffold similarity (Scaf), and Fréchet ChemNet Distance (FCD) Polykovskiy et al. (2020).
  • 3D Metrics: Atom stability, FCD, maximum mean discrepancy (MMD) Gretton et al. (2012) for the distributions of bond lengths, bond angles, and dihedral angles.
The details of these metrics are in Appendix A.1.1. We also report all the metrics for molecules from training set, serving as an approximate upper bound for performance.
Table 2 and Table 3 present the performance of UDM-3D on the de novo generation task on QM9 and GEOM-Drugs datasets, respectively. Our model demonstrates leading performances in generating chemically valid and novel molecules and achieves state-of-the-art performance in most of the metrics across both datasets. Two detailed key observations emerge from the results:
  1. 1.
    Breakthrough in Geometric Accuracy. Our unified latent space proves highly effective for 3D geometric modeling which is crucial for accurate 3D molecule generation. UDM-3D achieves significantly better performance than baselines for the distributional distance of bond lengths, bond angles, reducing errors by an order of magnitude. This demonstrates UDM-3D’s ability to generate highly accurate 3D structures. For the distributional distance of dihedral angles, a key measurement for 3D conformation accuracy Jing et al. (2022), UDM-3D also achieves the best performance. Notably, these advantages on stereochemical geometric structures persist for complex drug-like molecules in GEOM-Drugs, where UDM-3D improves the distributional distance for bond length by 25 times compared to GeoLDM (9.89E-03 v.s. 3.91E-01). This quantum leap stems from UDM-3D’s advantage to jointly optimize chemical and 3D geometric constraints when generating within a unified latent space, ensuring consistency across different molecular modalities.
  2. 2.
    Novel and High Quality Generation. UDM-3D achieves remarkable V&U&N scores of 0.948 (QM9) and 0.958 (GEOM-Drugs) while maintaining stability on par with top baselines like JODO. This demonstrates UDM-3D’s unique capability to generate novel molecules while preserving high-quality structures. The improvement is attributed to the unified latent space: by encoding atom types and bond features in a unified space, the model avoids overfitting to common molecular patterns while maintaining chemical validity.
4.3 Conditional 3D Molecule Generation
Experimental Settings. We evaluate conditional generation of molecules with target quantum properties using the protocol from Hoogeboom et al. (2022); Xu et al. (2023). Specifically, our target properties include $C_{v}$, $\mu$, $\alpha$, $\epsilon_{\text{HOMO}}$, $\epsilon_{\text{LUMO}}$, $\Delta\epsilon$, and we evaluate the Mean Absolute Error (MAE) between target and predicted properties. More details on the properties and settings are provided in Appendix A.1.2.
Table 4 presents the MAE results for conditional generation on QM9. UDM-3D achieves the lowest MAE across five properties, demonstrating its effectiveness in generating molecules with desired target properties. Notably, UDM-3D reduces the average MAE by 42.2% compared to GeoLDM, with strong gains (52.7% relative improvement) in $\Delta\epsilon$ (HOMO-LUMO gap), demonstrating significant improvements in modeling quantum chemical properties. These gains highlight the critical role of our unified latent space in harmonizing multi-modal information during conditional generation. While GeoLDM struggles to correlate all the modalities with target quantum properties due to its fragmented latent spaces, our unified latent representation inherently preserves the interplay between 3D geometry, bonds, and electronic characteristics, enabling precise modeling over molecule-property relationships.
4.4 Training & Sampling Efficiency
(a) Training time comparison.
(b) Sampling speed comparison.
Figure 3: Efficiency comparison. (a): Our UDM-3D adopts a two-stage pipeline (stacked bar: 14h VAE training + 38h LDM training), while baselines use end-to-end training.
We further analyze the training and sampling efficiency of UDM-3D compared to baselines on the QM9 dataset. These experiments are performed with an NVIDIA A100 GPU.
Training Efficiency. As shown in Figure 3(a), UDM-3D completes training in 52 hours (14h UAE-3D + 38h UDM-3D), which is 5.3 times faster than GeoLDM (449h) and 2.7 times faster than JODO (139h). This efficiency stems from: (1) Our decoupled training paradigm: UAE-3D first learns 3D molecule compression, allowing the DiT to focus solely on generative modeling of the compressed latents. (2) Efficient diffusion with DiT: Unlike previous molecular diffusion models that require complex architectures for multi-modality and equivariance, DiT’s simple and highly parallelizable design allows for faster training and sampling on the unified latent space. Crucially, our speedup does not compromise geometric accuracy, as guaranteed by UAE-3D’s near-lossless molecular reconstruction (cf. Table 1).
Sampling Speed. As shown in Figure 3(b), UDM-3D generates molecules in just 0.081s per sample, making it 7.3x faster than EDM/GeoLDM (0.59s) and 9.8x faster than JODO (0.79s). This speed advantage comes from the unified latent space, which simplifies DiT’s modeling and avoids complex neural architectures.
Table 5: Comparing DiT with vanilla Transformer for latent diffusion modeling. Models have the same depth and hidden size. A.S. stands for atom stability
Arch. A.S.${}_{\text{3D}}$ A.S.${}_{\text{2D}}$ V&C V&U V&U&N
DiT 0.993 0.999 0.983 0.972 0.948
Trans. 0.983 0.997 0.938 0.922 0.922
4.5 Ablation Study and Hyperparameter Analysis
DiT Backbone. We conduct an ablation study on DiT by comparing it to a vanilla Transformer for de novo 3D molecule generation on the QM9 dataset (Table 5). The results show that DiT outperforms the vanilla Transformer across all metrics, which we attribute to its adaptive LayerNorm layers. These layers enable DiT to effectively handle data with varying noise scales, improving diffusion performance.
5 Related Works
3D Molecule Generation. Early efforts for 3D molecule generation focus on autoregressive approaches Gebauer et al. (2019, 2021); Luo and Ji (2022), constructing 3D molecules sequentially by adding atoms or molecular fragments. With the success of diffusion models across various domains Sohl-Dickstein et al. (2015); Ho et al. (2020); Dhariwal and Nichol (2021), they also become the de facto method for 3D molecule generation Hoogeboom et al. (2022); Bao et al. (2023). However, the early diffusion works can easily generate invalid molecules because of overlooking the bond information. To bridge this gap, subsequent works Huang et al. (2023b); Vignac et al. (2023); Hua et al. (2023) additionally consider bond generation by building a separate diffusion process. However, generating different modalities in separate diffusion processes unnecessarily complicates the model design, reducing efficiency. It also risks compromising the consistency between these modalities. To address this, our UDM-3D performs generative modeling in UAE-3D’s unified latent space that integrates all molecular modalities, improving efficiency and generation quality.
Latent Diffusion Models (LDMs) for 3D Molecule Generation. To improve the efficiency, LDMs employ a VAE to compress raw data into a lower-dimensional latent space, where a diffusion model performs generative modeling Rombach et al. (2022). This approach has been very popular for image generation Peebles and Xie (2023). However, existing LDM-based methods for 3D molecule generation still face challenges in separated latent spaces. For example, GeoLDM Xu et al. (2023) compresses atomic features and 3D coordinates separately, failing to achieve a unified latent space distribution. In 3D protein molecule generation, PepGLAD Kong et al. (2024) compresses 1D sequences and 3D structures into two separate latent spaces. Our UDM-3D addresses these issues by employing a unified latent space, significantly improved generation efficiency.
Molecular Variational Autoencoders. Variational Autoencoders (VAEs) have been widely explored for generative modeling Kingma and Welling (2014); Simonovsky and Komodakis (2018); van den Oord et al. (2017); Razavi et al. (2019), providing a framework to encode data into a structured latent space while enabling both generation and reconstruction Kingma and Welling (2014). In molecular generation, VAEs have been applied to 2D molecular graphs, as seen in JT-VAEJin et al. (2018), MGCVAELee and Min (2022), SSVAEKang and Cho (2019), and CGVAE Liu et al. (2018). However, these methods have several limitations. First, they primarily focus on 2D molecular representations and often rely on non-transformer architectures, limiting their scalability for more complex molecular generation tasks Du et al. (2022). Additionally, VAE-based methods typically assume a simple Gaussian prior, which may fail to accurately model the complex posterior distributions required for effective molecular generation Dai and Wipf (2019); Aneja et al. (2021). With recent advancements in Latent Diffusion Models (LDMs) demonstrating success across multiple domains Vahdat et al. (2021); Zeng et al. (2022), LDMs have been introduced for 3D molecule and protein generation, as seen in GeoLDMXu et al. (2023) and PepGLADKong et al. (2024). These approaches leverage diffusion modeling in latent space, offering improved generative capacity and efficiency.
6 Conclusion and Future Works
In this paper, we propose UAE-3D to compress the multi-modal features of 3D molecules into a unified latent space, and demonstrate the effectiveness of latent diffusion modeling on this space by introducing UDM-3D. By integrating atom types, chemical bonds, and 3D coordinates into a single latent space, our model effectively addresses the inherent challenges of multi-modality and SE(3) equivariance for 3D molecule generation. Extensive experiments on GEOM-Drugs and QM9 confirm leading performances in both de novo and conditional 3D molecule generation, setting new benchmarks for both quality and efficiency.
The success of UAE-3D highlights the benefits of building a unified latent space for molecular design. Moving forward, we will investigate the transferability of the unified latent space to broader molecular design tasks, including structure-based drug design, and extend it to additional modalities, such as proteins, to further assess its effectiveness.
Appendix A Appendix
A.1 Experimental Details
A.1.1 Evaluation Metrics
Our evaluation framework for de novo 3D molecule generation encompasses complementary metrics to assess molecular validity, diversity, and geometric fidelity. These metrics operate at two levels:
2D Structural Analysis:
  • Validity & Stability: Measures adherence to chemical rules:
    • Atom Stability: Percentage of atoms with chemically valid valency
    • Validity & Completeness (V&C): Fraction of fully connected, syntactically correct molecules
  • Diversity: Quantifies exploration of chemical space:
    • V&U/V&U&N: Rates uniqueness and novelty relative to training data
  • Distribution Alignment: Compares generated/test distributions:
    • FCD: Fréchet distance in ChemNet feature space
    • SNN/Frag/Scaf: Tanimoto similarity for nearest neighbors, BRICS fragments, and Murcko scaffolds
3D Geometric Analysis:
  • Conformer Quality: Evaluates reconstructed 3D coordinates:
    • Atom Stability: Consistency of valency in 3D-derived 2D graphs
[…]
  • Geometric Fidelity: Measures spatial arrangement accuracy via Maximum Mean Discrepancy (MMD) for:
    • Bond lengths
    • Bond angles
    • Dihedral angles
A.1.2 Targeted Quantum Properties
For conditional generation, we optimize molecules toward six key electronic properties from quantum chemistry:
  • Electronic Response:
    • Dipole Moment ($\mu$): Molecular polarity measure
    • Polarizability ($\alpha$): Induced dipole response to electric fields
  • Thermodynamic Properties:
    • Heat Capacity ($C_{v}$): Energy absorption at constant volume
  • Electronic Structure:
    • HOMO ($\varepsilon_{\text{HOMO}}$)/LUMO ($\varepsilon_{\text{LUMO}}$): Frontier orbital energies
    • HOMO-LUMO Gap ($\Delta\varepsilon$): Critical for reactivity and conductivity
The evaluation protocol follows a rigorous split-and-validate strategy Hoogeboom et al. (2022):
  • QM9 training data divided into disjoint subsets $D_{a}$ (50k) and $D_{b}$ (50k)
  • Property predictor $\phi_{c}$ trained exclusively on $D_{a}$
  • Generated molecules compared against $\phi_{c}$’s predictions on $D_{b}$ (lower-bound baseline) and random predictions without any relation between molecule and property (upper-bound baseline)
  • Performance quantified via Mean Absolute Error (MAE) across all properties
This approach ensures fair assessment of conditional generation without data leakage, with $\phi_{c}$’s $D_{b}$ performance establishing the theoretical minimum achievable error.
A.2 Implement Details
Algorithm 1 Training Algorithm for UAE-3D and UDM-3D
0:  Molecular dataset $\mathcal{D}$, encoder $\mathcal{E}$, decoder $\mathcal{D}$, diffusion model $\epsilon_{\theta}$
1:  Initialize encoder $\mathcal{E}$, decoder $\mathcal{D}$, and diffusion model $\epsilon_{\theta}$
2:  Stage 1: Train UAE-3D
3:  while not converged do
4:     Sample a batch of 3D molecules $\mathbf{G}\sim\mathcal{D}$
5:     $\mathbf{Z}=\mathcal{E}(\mathbf{G})$Encode}
6:     $\hat{\mathbf{G}}=\mathcal{D}(\mathbf{Z})$Decode}
7:     $\mathcal{L}_{\text{recon}}$ (Eq. 20) {Reconstruction loss}
8:     $\mathcal{L}_{\text{UAE-3D}}$ (Eq. 21) {VAE loss}
9:     Update $\mathcal{E}$ and $\mathcal{D}$ using $\mathcal{L}_{\text{UAE-3D}}$
10:  end while
11:  Stage 2: Train UDM-3D
12:  while not converged do
13:     Sample a batch of latent sequences $\mathbf{Z}^{(0)}\sim\mathcal{E}(\mathcal{D})$
14:     Sample diffusion timestep $t\sim\text{Uniform}(0,1)$
15:     $\mathbf{Z}^{(t)}=\sqrt{\bar{\alpha}^{(t)}}\mathbf{Z}^{(0)}+\sqrt{1-\bar{\alpha}^{(t)}}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I})$Diffusion}
16:     $\hat{\boldsymbol{\epsilon}}=\epsilon_{\theta}(\mathbf{Z}^{(t)},t)$
17:     $\mathcal{L}_{\text{diffusion}}=\|\hat{\boldsymbol{\epsilon}}-\boldsymbol{\epsilon}\|^{2}$Noise prediction loss}
18:     Update $\epsilon_{\theta}$ using $\mathcal{L}_{\text{diffusion}}$
19:  end while
Algorithm 2 UDM-3D Sampling with CFG
0:  Trained UAE-3D encoder $\mathcal{E}$, DiT $\boldsymbol{\epsilon}_{\theta}$
0:  Guidance strength $w$, timesteps $\{t_{k}\}_{k=1}^{K}$
0:  Target property $c$ (optional)
1:  $\mathbf{Z}^{(1)}\sim\mathcal{N}(0,\mathbf{I})$Initial noise}
2:  for $k=K$ downto $1$ do
3:     $t\leftarrow t_{k}$
4:     Compute Guidance:
5:     $\boldsymbol{\epsilon}_{\text{cond}}\leftarrow\boldsymbol{\epsilon}_{\theta}(\mathbf{Z}^{(t)},t,c)$
6:     $\boldsymbol{\epsilon}_{\text{uncond}}\leftarrow\boldsymbol{\epsilon}_{\theta}(\mathbf{Z}^{(t)},t,\emptyset)$
7:     $\tilde{\boldsymbol{\epsilon}}\leftarrow(1+w)\boldsymbol{\epsilon}_{\text{cond}}-w\boldsymbol{\epsilon}_{\text{uncond}}$CFG blending}
8:     Denoise:
9:     $\mathbf{Z}^{(t-1)}\leftarrow\text{DDPM\_Step}(\mathbf{Z}^{(t)},\tilde{\boldsymbol{\epsilon}},t)$ Ho et al. (2020)
10:  end for
11:  Decode:
12:  $\hat{\mathbf{G}}=\langle\hat{\mathbf{F}},\hat{\mathbf{E}},\hat{\mathbf{X}}\rangle\leftarrow\mathcal{D}(\mathbf{Z}^{(0)})$UAE-3D decoder}
Hyperparameters. We refer insterested readers to our open-source code https://anonymous.4open.science/r/UAE-3D/ for a comprehensive view of our hyperparameters, such as learning rate and hidden dimension.
A.2.1 Algorithm
We provide the training and sampling algorithms for UDM-3D in Algorithms 1 and 2, respectively. The training algorithm for UDM-3D involves two stages: training the VAE (UAE-3D) and training the DiT (UDM-3D). The sampling algorithm for UDM-3D with CFG involves iteratively denoising the latent sequence $\mathbf{Z}$ using the DiT and the CFG guidance.